A multi-disc clutch has three discs on the driving shaft and two on the driven shaft. The inner diameter of the contact surface is 120 mm. The maximum pressure between the surface is limited to 0.1 N/mm2. Find the outer diameter for transmitting 25 kW at 1575 rpm. Assume uniform wear condition and coefficient of friction as 0.3.
1
Expert's answer
2020-06-05T05:36:58-0400

For uniform wear "p\cdot r=C=const" . Since


"p\cdot r_2=0.1\cdot120=12N/mm"


Axial force required to engage the clutch


"W=2\pi C(r_1-r_2)"


The torque transmitted


"T=4\cdot\mu W\frac{r_1+r_2}{2}=4\cdot \mu\cdot 2\pi C(r_1-r_2)\cdot\frac{r_1+r_2}{2}=4\cdot\mu\cdot \pi C(r_1-r_2)\cdot (r_1+r_2)"


Power transmitted


"P=T\cdot \omega=4\cdot\mu\cdot \pi C(r_1-r_2)\cdot (r_1+r_2)\cdot 2\pi \nu" "\to"


"r_1=\sqrt{\frac{P}{8\mu\pi^2C\nu}+r_2^2}=\sqrt{\frac{25000}{8\cdot0.3\cdot3.14^2\cdot12000\cdot26.25}+0.12^2}=0.158m=158mm" Answer.